Circular No. 6759

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS D. A. Biesecker, University of Birmingham, on behalf of the SOHO-

LASCO Consortium (cf. IAUC 6685), reports his discovery of two more comets, evidently Kreutz sungrazers, in C3 and C2 coronagraphic data. Each showed evidence for a small tail and was estimated to peak at mag ∼ 7. Measurements by Biesecker (assisted by C. St. Cyr for the second comet) have been reduced by G. V. Williams and are given in detail on MPEC 1997-U06 and -U07, together with orbital solutions by the undersigned.

1997 UT  α_{2000}  δ_{2000}
Oct. 6.218  12^{h}31.^{m}7  − 6^{o}06′  C/1997 T4
7.217  12 32.5  − 6 48  C/1997 T5

COMET P/1997 T3 G. Hahn, DLR Institute of Planetary Exploration, Berlin, commu-

nicates his precise measurement of the then-apparently-asteroidal object, which was discovered by Claes-Ingvar Lagerkvist, Uppsala University Observatory, on a film obtained with the European Southern Observatory’s 1-m Schmidt on Oct. 1 in the course of the Uppsala-DLR Trojan Survey. New orbital elements from the full Oct. 1–18 arc are given on MPEC 1997-U08. A Jupiter-Trojan orbit (cf. IAUC 6755) is no longer viable.

1997 UT  α_{2000}  δ_{2000}
Oct. 1.23681  1^{h}17^{m}11.^{s}78  +3^{o}13′44′′.0

SEVENTY-FIFTH ANNIVERSARY ISSUE The fiftieth anniversary of the first IAU Circular was noted on IAUC

2452 on 1972 Oct. 20. The present Circular marks the seventy-fifth anniversary, which has also been acknowledged by the naming of (7608) Telegramia, a minor planet, provisionally designated 1995 UO_{1}, discovered by J. Tichá at the Kleť Observatory on 1995 Oct. 22. The following citation for this naming is extracted from MPC 30802 (dated 1997 Oct. 16): “Named on the 75th anniversary of the publication of the first IAU Circular, which was issued on 1922 Oct. 22 by the IAU Central Bureau for Astronomical Telegrams, then at the Copenhagen Observatory. The Bureau moved to the Smithsonian Astrophysical Observatory in 1965 and has now issued 6756 IAU Circulars, relating generally to transient astronomical objects and phenonena in urgent need of further observation. Name proposed by the discoverer and endorsed by her colleagues M. Tichý and Z. Moravec.”

1997 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden